OpenStudy (itsmichelle29):

What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3? x + 3y = 10 x + 3y = −10 3x + y = 10 3x + y = −10

OpenStudy (itsmichelle29):

help plz

OpenStudy (itsmichelle29):

help please

OpenStudy (imtiaz7):

here we use point slope form that is y-y1=m(x-x1) here y1=-2 and x1=4 and m=-3 now just put the valuse you will get the answer

OpenStudy (itsmichelle29):

so ill get y - -2 = -3 x-4

OpenStudy (itsmichelle29):

then what

OpenStudy (imtiaz7):

wait plz

OpenStudy (imtiaz7):

C is your answer! :)

OpenStudy (itsmichelle29):

thank u so much

OpenStudy (imtiaz7):

welcome so itsmichelle :)